A stock originally purchased exactly 6 years ago for $231.00 per share is currently valued at $318.25 per share - what is the Geometric Mean Return generated by this investment since its original purchase date (assume annual compounding)?

-4.6%

6.3%

5.8%

-5.2%

5.2%

0.4%

5.5%
Answer #1

Original Purchase Price = P0 = $231.00

Current Purchase Price = P1 = $318.25

Let geometic return = r

n = 6 years

P1 = P0 *(1+r)^n

$318.25 = $231.00 * (1+r)^6

(1+r)^6 = 1.37770563

1+r = 1.05485494

r = 0.05485494

Therfore, geometric return is 5.50%